Consider the function f(x)=2x^2-0.9x-1 and calculate f (0) and f (1.5) in R program
Answer by stanbon(75887)

Consider the function f(x)=2x^2-0.9x-1 and calculate
f (0) = -1
and
f(1.5) = 2(1.5)^2 - 0.9(1.5) -1 = 2(2.25) - 1.35 - 1 = 4.5-2.35 = 2.15
-----------------------
Cheers,
Stan H.
----------------